OpenStudy (anonymous):

A farmer decided to fence a triangular vegetable garden from his field bounded on two sides by two streets intersecting at 13'15'. the sides of the garden along the streets are 10 meters and 5 meters. Find the perimeter of the garden. Find the area of the garden. Seedlings are to be planted in the garden at a cost of Php15.00 per foot. What's the total cost?

OpenStudy (confusionist):

1. The perimeter is: 10+5+11.18 =26.18

OpenStudy (confusionist):

You get that by adding all sides. It should be pretty easy. Then, multiply width by length to get the triangle area. Be sure to divide this by two. 10*5= 50, 25

OpenStudy (confusionist):

For the last one all you have to do is multiply 15*25=375.